# Decision making in stochastic extensive form I: Stochastic decision forests

### v1

@@ 2.2 Stochastic decision forests
Let $(F,\pi,{\bm{\mathsf{X}}})$ be a stochastic decision forest over some exogenous scenario space $(\Omega,\mathscr{A})$ such that all roots are moves. [^9] Of course, any sdf can be reduced to this form by eliminating the connected components (alias exogenous scenarios) without moves. Then, $({\bm{\mathsf{T}}},\geq_{\bm{\mathsf{T}}})$ defines a rooted decision tree and ${\bm{\mathsf{X}}}$ is the set of its moves.
How does an agent (alias decision maker) pass through a stochastic decision forest? There are actually two stories that can be told. On the one hand, as motivated in the outset of this subsection, first an exogenous scenario $\omega\in\Omega$ is determined, and then the agent passes through the decision tree given by the connected component $\pi^{-1}(\{\omega\})$. On the other hand, the agent passes through the rooted decision tree $({\bm{\mathsf{T}}},\geq_{\bm{\mathsf{T}}})$. While the first perspective is the correct one in respect to outcome generation, the second one, though, is more relevant for the description of the exogenous information available to agents. The agents may not know $\omega$ or, even if they do, that knowledge may not be determinant. [^10] That point will be discussed in detail in subsection 3.1. The point of random moves lies in the fact that they serve as a flexible and general basis for “oracles”, that is, a structure of exogenous information available to an agent. They will form the points that exogenous information is revealed at, and it is the tree property of $({\bm{\mathsf{T}}},\geq_{\bm{\mathsf{T}}})$ that will retain the extensive-form character of that process. In turn, the order-theoretic properties of a random move themselves do not reveal any exogenous information – this is a big difference to the classical approach to stochastic games and decision problems in extensive form, where exogenous information is the result of a “nature” player or agent making dynamic choices (going back to Shapley’s work [39], see, e.g., [5, section 2.2.2.5]).
Note that stochastic decision forests are allowed to vary across scenarios: the trees in $\mathcal{F}$ need not be isomorphic. The crowns may become shallower in some scenarios (because some options are no more available there).
Remark 2.13.
It can be checked in the proof in the appendix that proposition 2.11 holds true even if we do not assume properties 3d, 3e, and 3f.
@@ 2.3 Simple examples
In the following we present two very simple examples of stochastic decision forests. The first example is illustrated in figure 1.
${\bm{\mathsf{x}}}_{0}(\omega_{1})$${\bm{\mathsf{x}}}_{1}(\omega_{1})$${\bm{\mathsf{x}}}_{2}(\omega_{1})$${\bm{\mathsf{x}}}_{1}(\omega_{2})$${\bm{\mathsf{x}}}_{0}(\omega_{2})$${\bm{\mathsf{x}}}_{2}(\omega_{2})$$w_{112}$$w_{111}$$w_{122}$$w_{121}$$w_{211}$$w_{212}$$w_{221}$$w_{222}$
Figure 1: A simple stochastic decision forest represented as a directed graph, with $w_{\omega km}=(\omega,k,m)$, for $(\omega,k,m)\in W$; moves are indicated by circles
@@
Let $\omega_{1}=1$, $\omega_{2}=2$ and consider $\Omega=\{\omega_{1},\omega_{2}\}$, $\mathscr{A}=\mathcal{P}(\Omega)$, $W=\Omega\times\{1,2\}^{2}$ and ${\bm{\mathsf{x}}}_{0},{\bm{\mathsf{x}}}_{1},{\bm{\mathsf{x}}}_{2}\colon\Omega\to\mathcal{P}(W)$ given by ${\bm{\mathsf{x}}}_{0}(\omega)=\{\omega\}\times\{1,2\}^{2}$ and ${\bm{\mathsf{x}}}_{k}(\omega)=\{(\omega,k)\}\times\{1,2\}$, $F=\{{\bm{\mathsf{x}}}_{k}(\omega)\mid\omega\in\Omega,\leavevmode\nobreak\ k=0,1,2\}\cup\{\{w\}\mid w\in W\}$, $\pi\colon F\to\Omega$ be the map sending any node to the first entry of an arbitrary choice among its elements.
@@
The tuple $(F,\pi,{\bm{\mathsf{X}}})$ defines a stochastic decision forest.
@@
${\bm{\mathsf{x}}}_{0}$${\bm{\mathsf{x}}}_{1}$${\bm{\mathsf{x}}}_{2}$$w_{121}$$w_{112}$$w_{111}$$w_{122}$$w_{212}$$w_{221}$$w_{211}$$w_{222}$
Figure 2: The decision tree $({\bm{\mathsf{T}}},\geq_{\bm{\mathsf{T}}})$ for the simple stochastic decision forest, with $w_{\omega km}=(\omega,k,m)$, for $(\omega,k,m)\in W$; (random) moves are indicated by circles
As a variant, identifying the elements $(\omega_{1},2,1)$ and $(\omega_{1},2,2)$ in $W$ will provide a stochastic decision forest with a random move that is not defined on all of $\Omega$, as illustrated in figure 3.
${\bm{\mathsf{x}}}^{\prime}_{0}(\omega_{1})$${\bm{\mathsf{x}}}^{\prime}_{1}(\omega_{1})$$w^{\prime}_{12}$${\bm{\mathsf{x}}}^{\prime}_{0}(\omega_{2})$${\bm{\mathsf{x}}}^{\prime}_{1}(\omega_{2})$${\bm{\mathsf{x}}}^{\prime}_{2}(\omega_{2})$$w^{\prime}_{112}$$w^{\prime}_{111}$$w^{\prime}_{211}$$w^{\prime}_{212}$$w^{\prime}_{221}$$w^{\prime}_{222}$
Figure 3: A variant of the simple stochastic decision forest in figure 1 represented as a directed graph, with $w^{\prime}_{\omega km}=(\omega,k,m)$, for all triples $(\omega,k,m)\in W^{\prime}$, and $w^{\prime}_{12}=(\omega_{1},2)=(1,2)$; moves are indicated by circles
Put differently, let $W^{\prime}=W\setminus\{(\omega_{1},2,1),(\omega_{1},2,2)\}\cup\{(\omega_{1},2)\}$ and let ${\bm{\mathsf{x}}}^{\prime}_{0}={\bm{\mathsf{x}}}_{0}$, ${\bm{\mathsf{x}}}^{\prime}_{1}={\bm{\mathsf{x}}}_{1}$, and ${\bm{\mathsf{x}}}^{\prime}_{2}\colon\{\omega_{2}\}\to\mathcal{P}(W^{\prime})$ be given by ${\bm{\mathsf{x}}}^{\prime}_{2}(\omega_{2})=\{(\omega_{2},2)\}\times\{1,2\}$. Let ${\bm{\mathsf{X}}}^{\prime}=\{{\bm{\mathsf{x}}}^{\prime}_{0},{\bm{\mathsf{x}}}^{\prime}_{1},{\bm{\mathsf{x}}}^{\prime}_{2}\}$. Let $D_{{\bm{\mathsf{x}}}^{\prime}_{0}}=\Omega$, $D_{{\bm{\mathsf{x}}}^{\prime}_{1}}=\Omega$, $D_{{\bm{\mathsf{x}}}^{\prime}_{2}}=\{\omega_{2}\}$ and $F^{\prime}=\{{\bm{\mathsf{x}}}^{\prime}(\omega)\mid{\bm{\mathsf{x}}}^{\prime}\in{\bm{\mathsf{X}}}^{\prime},\leavevmode\nobreak\ \omega\in D_{{\bm{\mathsf{x}}}^{\prime}}\}\cup\{\{w^{\prime}\}\mid w^{\prime}\in W^{\prime}\}$. Let $\pi\colon F^{\prime}\to\Omega$ be the map sending any node to the first entry of an arbitrary choice among its elements.
@@
The tuple $(F^{\prime},\pi^{\prime},{\bm{\mathsf{X}}}^{\prime})$ defines a stochastic decision forest.
${\bm{\mathsf{x}}}^{\prime}_{0}$${\bm{\mathsf{x}}}^{\prime}_{1}$$w^{\prime}_{12}$${\bm{\mathsf{x}}}^{\prime}_{2}$$w^{\prime}_{112}$$w^{\prime}_{111}$$w^{\prime}_{211}$$w^{\prime}_{212}$$w^{\prime}_{221}$$w^{\prime}_{222}$
Figure 4: The decision tree $({\bm{\mathsf{T}}},\geq_{\bm{\mathsf{T}}})$ for the variant of the simple stochastic decision forest, with $w^{\prime}_{\omega km}=(\omega,k,m)$, for all triples $(\omega,k,m)\in W^{\prime}$, and $w^{\prime}_{12}=(\omega_{1},2)=(1,2)$; (random) moves are indicated by circles
@@ 2.4 Action path stochastic decision forests
In most pieces of the literature, dynamic games are defined by supposing a notion of time and specifying outcomes as certain paths of action at instants of time. [3, Subsection 2.2] provides a broad overview for this, including classical textbook definitions as in [21], infinite bilateral bargaining in discrete time as in [37], repeated games, the long cheap talk game in [7], and a decision-theoretic interpretation of differential games as in [18]. We desire to add, in a first approximate sense, stochastic control in both discrete and continuous time (see, e.g., [34, 10, 29]) and stochastic differential games (see, e.g., [13]) without restrictions on the noise in question.
In this subsection, this approach is formulated in one abstract and general framework. This framework is based on a specific structure pertaining to all of these examples, namely time. Interestingly, time is not included in the abstract formulation of decision forests, and it serves as a particularly strong similarity structure for trees and even branches of one and the same tree. We do not address the question whether all (stochastic) decision forests can be represented in this framework. Of course, the second major point of the following framework is that it will allow for general exogenous stochastic noise, going strictly beyond the “nature” agent setting.
Let $\mathbb{T}\subseteq\mathbb{R}_{+}$ be such that $0\in\mathbb{T}$ and $\mathbb{A}$ be a non-empty metric space. Let $(\Omega,\mathscr{A})$ be an exogenous scenario space. Let $W\subseteq\Omega\times\mathbb{A}^{\mathbb{T}}$ be such that for all $\omega\in\Omega$, there is $f\in\mathbb{A}^{\mathbb{T}}$ with $(\omega,f)\in W$. An outcome will thus be a pair of an exogenous scenario and a path $f\colon\mathbb{T}\to\mathbb{A}$ of “action”, and any scenario is required to admit at least one outcome. For any $t\in\mathbb{T}$ and $\tilde{w}=(\omega,f)\in\Omega\times\mathbb{A}^{\mathbb{T}}$, let
$x_{t}(\tilde{w})=x_{t}(\omega,f)=\{(\omega^{\prime},f^{\prime})\in W\mid\omega^{\prime}=\omega,\leavevmode\nobreak\ f^{\prime}|_{[0,t)_{\mathbb{T}}}=f|_{[0,t)_{\mathbb{T}}}\}.$
Let $F=\{x_{t}(w)\mid t\in\mathbb{T},\leavevmode\nobreak\ w\in W\}\cup\{\{w\}\mid w\in W\}$. Further, let $\pi\colon F\to\Omega$ be the unique function mapping any $x\in F$ to the first item of one choice of its elements.
@@
  • 1.
    Assumption AP.W0. For all $t\in\mathbb{T}$ and $f\in\mathbb{A}^{\mathbb{T}}$, $D_{t,f}\in\mathscr{A}$.
  • 2.
    Assumption AP.W1. For all $w\in W$ and all $t,u\in\mathbb{T}$ with $t\neq u$ and $x_{t}(w)=x_{u}(w)$, we have $x_{t}(w)=\{w\}$.
  • 3.
    Assumption AP.W2. For all $\omega\in\Omega$ and $\tilde{f}\in\mathbb{A}^{\mathbb{T}}$, and for all subsets $\mathbb{T}^{\prime}\subseteq\mathbb{T}$ satisfying $x_{t}(\omega,\tilde{f})\in F$ for all $t\in\mathbb{T}^{\prime}$, there is $f\in\mathbb{A}^{\mathbb{T}}$ with $(\omega,f)\in W$ and $f|_{[0,t)_{\mathbb{T}}}=\tilde{f}|_{[0,t)_{\mathbb{T}}}$ for all $t\in\mathbb{T}^{\prime}$.
  • 4.
    Assumption AP.W3. For all $t\in\mathbb{T}$ and $f,g\in\mathbb{A}^{\mathbb{T}}$ such that $D_{t,f},D_{t,g}\neq\emptyset$ and $D_{t,f}\cap D_{t,g}=\emptyset$, there is $u\in[0,t)_{\mathbb{T}}$ such that $D_{u,f}\cap D_{u,g}\neq\emptyset$ and $f|_{[0,u)_{\mathbb{T}}}\neq g|_{[0,u)_{\mathbb{T}}}$.
  • 5.
    Assumption AP.W4. There is a non-empty finite set $I$ and a family of metric spaces $\mathbb{A}^{i}$, $i\in I$, such that $\mathbb{A}=\prod_{i\in I}\mathbb{A}^{i}$ in the category of topological spaces.
Assumption AP.W0 requires that $D_{t,f}$ is indeed an event in $(\Omega,\mathscr{A})$. AP.W1 stipulates that any (presumptive) move has a unique time associated to it. Indeed, we have the following result. For any $x\in F$, let
@@
Assumption AP.W1 is satisfied iff for all $x\in F$ that are not singletons, $\mathbb{T}_{x}$ is a singleton.
Note that the following weak converse statement holds true independently of whether AP.W1 is assumed or not:
@@
If assumption AP.W1 is satisfied and $\mathbb{T}$ has no maximum, then for all $x\in F$ the following statements are equivalent:
  • 1.
@@
Assumption AP.W2 corresponds to what is generally called “boundedness” in [3, subsection 3.4]. In the present context, this is the conditio sine qua non demanding that any (presumptive) decision path, alias maximal chain of nodes, corresponds to a path of action $\mathbb{T}\to\mathbb{A}$. Assumption AP.W3 refers to the maximality axiom 3e and ensures that different paths in different scenarios cannot be identified. If the assumption does not hold true, then it may be possible to identify the path $f$ on $D_{t,f}$ with the path $g$ on $D_{t,g}$, as for all $u<t$ until that (exclusively) $f$ and $g$ can be distinguished, $D_{u,f}$ and $D_{u,g}$ continue to be disjoint. Finally, assumption AP.W4 describes the situation where an action is the product of the individual actions of a finite set of $I$ “agents”. Of course, this assumption is trivially satisfied for singleton $I$ and $\mathbb{A}^{i}=\mathbb{A}$ for the unique $i\in I$, but sometimes it is satisfied for larger $I$. Hence, the relevant point about AP.W4 is not whether it is satisfied, but with respect to what family $(\mathbb{A}^{i})_{i\in I}$ it is.

### v2

@@ 2.2 Stochastic decision forests
Let $(F,\pi,{\bm{\mathsf{X}}})$ be an order consistent, surely non-trivial and maximal stochastic decision forest on some exogenous scenario space $(\Omega,\mathscr{E})$. Then, $({\bm{\mathsf{T}}},\geq_{\bm{\mathsf{T}}})$ defines a rooted decision tree and ${\bm{\mathsf{X}}}$ is the set of its moves.
Note, however, that $F$ cannot be subsumed under this derived object. The latter as such does not faithfully account for the fact that there is a realised scenario. For instance, only those maximal chains $\mathbf{c}$ in $({\bm{\mathsf{T}}},\geq_{\bm{\mathsf{T}}})$ correspond to outcomes in $W=\bigcup F$ that satisfy $\bigcap_{{\bm{\mathsf{y}}}\in\mathbf{c}}D_{\bm{\mathsf{y}}}\neq\emptyset$, which need not be the case. This is developed a bit further in the second part, [40, Subsection 3.1].
@@ 2.3 Simple examples
In the following we present three very simple examples of stochastic decision forests. The first example is illustrated in Figure 1.
${\bm{\mathsf{x}}}_{0}(\omega_{1})$${\bm{\mathsf{x}}}_{1}(\omega_{1})$${\bm{\mathsf{x}}}_{2}(\omega_{1})$${\bm{\mathsf{x}}}_{1}(\omega_{2})$${\bm{\mathsf{x}}}_{0}(\omega_{2})$${\bm{\mathsf{x}}}_{2}(\omega_{2})$$\{w_{112}\}$$\{w_{111}\}$$\{w_{122}\}$$\{w_{121}\}$$\{w_{211}\}$$\{w_{212}\}$$\{w_{221}\}$$\{w_{222}\}$
Figure 1: A simple stochastic decision forest represented as a directed graph, with $w_{\ell km}=(\omega_{\ell},k,m)$, for $(\ell,k,m)\in\{1,2\}^{3}$. Moves are indicated by circles.
@@
Let $\Omega=\{\omega_{1},\omega_{2}\}$ be some exogenous scenario space with two elements and $\mathscr{E}=\mathcal{P}\Omega$. Let $W=\Omega\times\{1,2\}^{2}$ and ${\bm{\mathsf{x}}}_{0},{\bm{\mathsf{x}}}_{1},{\bm{\mathsf{x}}}_{2}\colon\Omega\to\mathcal{P}(W)$ given by ${\bm{\mathsf{x}}}_{0}(\omega)=\{\omega\}\times\{1,2\}^{2}$ and ${\bm{\mathsf{x}}}_{k}(\omega)=\{(\omega,k)\}\times\{1,2\}$, $F=\{{\bm{\mathsf{x}}}_{k}(\omega)\mid\omega\in\Omega,\leavevmode\nobreak\ k=0,1,2\}\cup\{\{w\}\mid w\in W\}$, $\pi\colon F\to\Omega$ be the map sending any node to the first entry of an arbitrary choice among its elements.
@@
The tuple $(F,\pi,{\bm{\mathsf{X}}})$ defines an order consistent, surely non-trivial, and maximal stochastic decision forest.
@@
${\bm{\mathsf{x}}}_{0}$${\bm{\mathsf{x}}}_{1}$${\bm{\mathsf{x}}}_{2}$$\{w_{212}\}_{\{\omega_{2}\}}$$\{w_{112}\}_{\{\omega_{1}\}}$$\{w_{111}\}_{\{\omega_{1}\}}$$\{w_{211}\}_{\{\omega_{2}\}}$$\{w_{122}\}_{\{\omega_{1}\}}$$\{w_{221}\}_{\{\omega_{2}\}}$$\{w_{121}\}_{\{\omega_{1}\}}$$\{w_{222}\}_{\{\omega_{2}\}}$
Figure 2: The decision tree $({\bm{\mathsf{T}}},\geq_{\bm{\mathsf{T}}})$ for the simple stochastic decision forest, with $w_{\ell km}=(\omega_{\ell},k,m)$, for $(\ell,k,m)\in\{1,2\}^{3}$. (Random) moves are indicated by circles. Elements of ${\bm{\mathsf{T}}}\setminus{\bm{\mathsf{X}}}$, of the form $\{(\omega,\{w\})\}$ and seen as maps $\omega\mapsto\{w\}$, are denoted by $\{w\}_{\{\omega\}}$.
As a variant, identifying the elements $(\omega_{1},2,1)$ and $(\omega_{1},2,2)$ in $W$ is seen below to provide a stochastic decision forest with a random move that is not defined on all of $\Omega$, as illustrated in Figure 3.
${\bm{\mathsf{x}}}^{\prime}_{0}(\omega_{1})$${\bm{\mathsf{x}}}^{\prime}_{1}(\omega_{1})$$\{w^{\prime}_{12}\}$${\bm{\mathsf{x}}}^{\prime}_{0}(\omega_{2})$${\bm{\mathsf{x}}}^{\prime}_{1}(\omega_{2})$${\bm{\mathsf{x}}}^{\prime}_{2}(\omega_{2})$$\{w^{\prime}_{112}\}$$\{w^{\prime}_{111}\}$$\{w^{\prime}_{211}\}$$\{w^{\prime}_{212}\}$$\{w^{\prime}_{221}\}$$\{w^{\prime}_{222}\}$
Figure 3: A variant of the simple stochastic decision forest in Figure 1 represented as a directed graph, with $w^{\prime}_{\ell km}=(\omega_{\ell},k,m)$, for all triples $(\ell,k,m)\in\{1,2\}^{3}$ with $(\omega_{\ell},k,m)\in W^{\prime}$, and $w^{\prime}_{12}=(\omega_{1},2)$. Moves are indicated by circles.
Put differently, let $W^{\prime}=W\setminus\{(\omega_{1},2,1),(\omega_{1},2,2)\}\cup\{(\omega_{1},2)\}$ and let ${\bm{\mathsf{x}}}^{\prime}_{0}={\bm{\mathsf{x}}}_{0}$, ${\bm{\mathsf{x}}}^{\prime}_{1}={\bm{\mathsf{x}}}_{1}$, and ${\bm{\mathsf{x}}}^{\prime}_{2}\colon\rho^{-1}(\{2\})\to\mathcal{P}(W^{\prime})$ be given by ${\bm{\mathsf{x}}}^{\prime}_{2}(\omega)=\{(\omega,2)\}\times\{1,2\}$. Let ${\bm{\mathsf{X}}}^{\prime}=\{{\bm{\mathsf{x}}}^{\prime}_{0},{\bm{\mathsf{x}}}^{\prime}_{1},{\bm{\mathsf{x}}}^{\prime}_{2}\}$. Let $D_{{\bm{\mathsf{x}}}^{\prime}_{0}}=\Omega$, $D_{{\bm{\mathsf{x}}}^{\prime}_{1}}=\Omega$, $D_{{\bm{\mathsf{x}}}^{\prime}_{2}}=\{\omega_{2}\}$ and $F^{\prime}=\{{\bm{\mathsf{x}}}^{\prime}(\omega)\mid{\bm{\mathsf{x}}}^{\prime}\in{\bm{\mathsf{X}}}^{\prime},\leavevmode\nobreak\ \omega\in D_{{\bm{\mathsf{x}}}^{\prime}}\}\cup\{\{w^{\prime}\}\mid w^{\prime}\in W^{\prime}\}$. Let $\pi^{\prime}\colon F^{\prime}\to\Omega$ be the map sending any node to the first entry of an arbitrary choice among its elements.
@@
The tuple $(F^{\prime},\pi^{\prime},{\bm{\mathsf{X}}}^{\prime})$ defines an order consistent, surely non-trivial, and maximal stochastic decision forest.
The corresponding decision tree $({\bm{\mathsf{T}}}^{\prime},\geq_{{\bm{\mathsf{T}}}^{\prime}})$ is illustrated in Figure 4.
${\bm{\mathsf{x}}}^{\prime}_{0}$${\bm{\mathsf{x}}}^{\prime}_{1}$$\{w^{\prime}_{12}\}_{\{\omega_{1}\}}$${\bm{\mathsf{x}}}^{\prime}_{2}$$\{w^{\prime}_{112}\}_{\{\omega_{1}\}}$$\{w^{\prime}_{111}\}_{\{\omega_{1}\}}$$\{w^{\prime}_{211}\}_{\{\omega_{2}\}}$$\{w^{\prime}_{212}\}_{\{\omega_{2}\}}$$\{w^{\prime}_{221}\}_{\{\omega_{2}\}}$$\{w^{\prime}_{222}\}_{\{\omega_{2}\}}$
Figure 4: The decision tree $({\bm{\mathsf{T}}}^{\prime},\geq_{{\bm{\mathsf{T}}}^{\prime}})$ for the variant of the simple stochastic decision forest, with $w^{\prime}_{\ell km}=(\omega_{\ell},k,m)$, for all triples $(\ell,k,m)\in\{1,2\}^{3}$ with $(\omega_{\ell},k,m)\in W^{\prime}$, and $w^{\prime}_{12}=(\omega_{1},2)$. (Random) moves are indicated by circles. Elements of ${\bm{\mathsf{T}}}^{\prime}\setminus{\bm{\mathsf{X}}}^{\prime}$, of the form $\{(\omega,\{w^{\prime}\})\}$ and seen as maps $\omega\mapsto\{w^{\prime}\}$, are denoted by $\{w^{\prime}\}_{\{\omega\}}$.
The third example is an sdf representation of Gilboa’s model of the absent-minded driver phenomenon in [22]. For this, let $(\Omega,\mathscr{E})$ be some exogenous scenario space and $\rho\colon\Omega\to\{1,2\}$ an $\mathscr{E}$-$\mathcal{P}\{1,2\}$-measurable surjection. Further, let $D,H,M$ be three pairwise distinct symbols meaning “disastrous region”, “home”, and “motel” as in the original story from [39]. Let $W=\Omega\times\{D,H,M\}$ and ${\bm{\mathsf{x}}}_{1},{\bm{\mathsf{x}}}_{2}$ be maps defined on the whole of $\Omega$ given by
${\bm{\mathsf{x}}}_{k}(\omega)=\begin{cases}\{\omega\}\times\{D,H,M\},&\text{ if }\rho(\omega)=k,\\ \{\omega\}\times\{H,M\},&\text{ if }\rho(\omega)=3-k,\end{cases}\qquad k\in\{1,2\}.$
$F$ is defined as the union of the images of ${\bm{\mathsf{x}}}_{1}$ and ${\bm{\mathsf{x}}}_{2}$ and the set of all singleton sets in $W$. $\pi\colon F\to\Omega$ maps any element of $F$ to the first component of its elements. $(F,\supseteq)$ and $({\bm{\mathsf{T}}},\geq_{\bm{\mathsf{T}}})$ are illustrated in Figure 5. Note that $({\bm{\mathsf{T}}},\geq_{\bm{\mathsf{T}}})$ is not even a forest, both ${\bm{\mathsf{x}}}_{1}$ and ${\bm{\mathsf{x}}}_{2}$ are maximal elements, and both $\operatorname{im}{\bm{\mathsf{x}}}_{1}$ and $\operatorname{im}{\bm{\mathsf{x}}}_{2}$ contain both a root and a move that is not a root, respectively.
${\bm{\mathsf{x}}}_{1}(\omega_{1})$$\{w_{1D}\}$${\bm{\mathsf{x}}}_{2}(\omega_{1})$$\{w_{1H}\}$$\{w_{1M}\}$${\bm{\mathsf{x}}}_{2}(\omega_{2})$$\{w_{1D}\}$${\bm{\mathsf{x}}}_{1}(\omega_{2})$$\{w_{1H}\}$$\{w_{1M}\}$
${\bm{\mathsf{x}}}_{1}$$\{w_{1D}\}_{\{\omega_{1}\}}$$\{w_{1H}\}_{\{\omega_{1}\}}$$\{w_{1M}\}_{\{\omega_{1}\}}$${\bm{\mathsf{x}}}_{2}$$\{w_{2D}\}_{\{\omega_{2}\}}$$\{w_{2H}\}_{\{\omega_{2}\}}$$\{w_{2M}\}_{\{\omega_{2}\}}$
Figure 5: The absent-minded driver sdf, following Gilboa: $(F,\supseteq)$ represented as a directed graph, in case $\rho$ is injective, with $\omega_{\ell}=\rho^{-1}(\ell)$, $w_{\ell S}=(\omega_{\ell},S)$, for all $\ell\in\{1,2\}$ and symbols $S$ such that $(\omega,S)\in W$ (left), $({\bm{\mathsf{T}}},\geq_{\bm{\mathsf{T}}})$ where elements of ${\bm{\mathsf{T}}}\setminus{\bm{\mathsf{X}}}$, of the form $\{(\omega,\{w\})\}$ and seen as maps $\omega\mapsto\{w\}$, are denoted by $\{w\}_{\{\omega\}}$ (right). Non-minimal elements are indicated by circles, respectively.
Lemma 2.16.
The tuple $(F,\pi,{\bm{\mathsf{X}}})$ just defined is a not order consistent stochastic decision forest.
@@ 2.4 Action path stochastic decision forests
In most pieces of the literature, dynamic games are defined by supposing a notion of time and specifying outcomes as certain paths of action at instants of time. [3, Subsection 2.2] provides a broad overview for this, including classical textbook definitions as in [21], infinite bilateral bargaining in discrete time as in [44], repeated games, the long cheap talk game in [7], and a decision-theoretic interpretation of differential games as in [18]. We desire to add to this list stochastic control in both discrete and continuous time (see, e.g. [38, 10, 33]) and stochastic differential games (see, e.g. [13]) without restrictions on the noise in question, and the first concrete step in this direction is taken in what follows next. [^9] As described in the introduction, this is a guiding theme of the whole series and its investigation culminates only in the third and final paper.
In this subsection, this approach is formulated in one abstract and general framework. This framework is based on a specific structure pertaining to all of these examples, namely time. Interestingly, time is not included in the abstract formulation of decision forests, and it serves as a particularly strong similarity structure for trees and even branches of one and the same tree. Of course, every decision path is totally ordered and therefore has its own time axis – but in the following construction, the “clocks” of all decision paths are be synchronised. [^10] We do not address the question whether all “relevant” order consistent and maximal stochastic decision forests can be represented in this framework – that is, admit a “clock” synchronising all their decision paths – though it does not seem obvious to conceive a counterexample. The second major point of the following framework is that it will allow for general exogenous stochastic noise, going strictly beyond the “nature” agent setting.
Let $(\mathbb{T},\leq)$ be a total order admitting a minimum which we denote by 00. Further, let $I$ be a non-empty finite set (which will represent decision makers alias agents, $(\mathbb{A}^{i})_{i\in I}$ be a family of non-empty metric spaces, and $\mathbb{A}=\prod_{i\in I}\mathbb{A}^{i}$ be their topological product, with canonical projections $p^{i}\colon\mathbb{A}\to\mathbb{A}^{i}$, $i\in I$. Of course, the case of singleton $I$ is included in this setting.
Let $(\Omega,\mathscr{E})$ be an exogenous scenario space. Let $W\subseteq\Omega\times\mathbb{A}^{\mathbb{T}}$ be such that for all $\omega\in\Omega$, there is $f\in\mathbb{A}^{\mathbb{T}}$ with $(\omega,f)\in W$. An outcome will thus be a pair of an exogenous scenario and a path $f\colon\mathbb{T}\to\mathbb{A}$ of “action”, and any scenario is required to admit at least one outcome. For any $t\in\mathbb{T}$ and $\tilde{w}=(\omega,f)\in\Omega\times\mathbb{A}^{\mathbb{T}}$, let
$x_{t}(\tilde{w})=x_{t}(\omega,f)=\{(\omega^{\prime},f^{\prime})\in W\mid\omega^{\prime}=\omega,\,f^{\prime}|_{[0,t)_{\mathbb{T}}}=f|_{[0,t)_{\mathbb{T}}}\}.$
Let $F=\{x_{t}(w)\mid t\in\mathbb{T},\,w\in W\}\cup\{\{w\}\mid w\in W\}$. Further, let $\pi\colon F\to\Omega$ be the unique function mapping any $x\in F$ to the first item of one choice of its elements.
@@
  • Assumption AP.SDF0. For all $t\in\mathbb{T}$ and $f\in\mathbb{A}^{\mathbb{T}}$, $D_{t,f}\in\mathscr{E}$.
  • Assumption AP.SDF1. For all $w\in W$ and all $t,u\in\mathbb{T}$ with $t\neq u$ and $x_{t}(w)=x_{u}(w)$, we have $x_{t}(w)=\{w\}$.
  • Assumption AP.SDF2. For all $\omega\in\Omega$ and $\tilde{f}\in\mathbb{A}^{\mathbb{T}}$, and for all subsets $\mathbb{T}^{\prime}\subseteq\mathbb{T}$ satisfying $x_{t}(\omega,\tilde{f})\in F$ for all $t\in\mathbb{T}^{\prime}$, there is $f\in\mathbb{A}^{\mathbb{T}}$ with $(\omega,f)\in W$ and $f|_{[0,t)_{\mathbb{T}}}=\tilde{f}|_{[0,t)_{\mathbb{T}}}$ for all $t\in\mathbb{T}^{\prime}$.
  • Assumption AP.SDF3. For all $t\in\mathbb{T}$ and $f,g\in\mathbb{A}^{\mathbb{T}}$ such that $D_{t,f},D_{t,g}\neq\emptyset$ and $D_{t,f}\cap D_{t,g}=\emptyset$, there is $u\in[0,t)_{\mathbb{T}}$ such that $D_{u,f}\cap D_{u,g}\neq\emptyset$ and $f|_{[0,u)_{\mathbb{T}}}\neq g|_{[0,u)_{\mathbb{T}}}$.
Assumption AP.SDF0 requires that $D_{t,f}$ is indeed an event in $(\Omega,\mathscr{E})$. Assumption AP.SDF1 stipulates that any (presumptive) move has a unique time associated to it. Indeed, we have the following result. For any $x\in F$, let
@@
Assumption AP.SDF1 is satisfied iff for all $x\in F$ that are not singletons, $\mathbb{T}_{x}$ is a singleton.
Note that the following weak converse statement holds true independently of whether AP.SDF1 is assumed or not:
@@
If Assumption AP.SDF1 is satisfied and $\mathbb{T}$ has no maximum, then for all $x\in F$ the following statements are equivalent:
@@
Assumption AP.SDF2 corresponds to what is generally called “boundedness” in [3, Subsection 3.4]. In the present context, this is the conditio sine qua non demanding that any (presumptive) decision path, alias maximal chain of nodes, corresponds to a path of action $\mathbb{T}\to\mathbb{A}$. Assumption AP.SDF3 refers to the maximality in Axiom 4(c) and ensures that different paths in different scenarios cannot be identified without violating the order consistency. If the assumption does not hold true, then it will become possible to identify the path $f$ on $D_{t,f}$ with the path $g$ on $D_{t,g}$, as for all $u<t$ until that (exclusively) $f$ and $g$ can be distinguished, $D_{u,f}$ and $D_{u,g}$ continue to be disjoint.